• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name nmuangma

    Log entry time 21:53:03 on June15,2010

    Entry number 326816

    keyword=target vacuum/ beam current

    at 21:17 the target vacuum fluctuate at higher pressure. Vincent decides to lower the beam and let the vacuum stable for awhile then we are now 21:42 back to the 50 micro-A.

    the vacuum yet stable for about 20 mins.


    FIGURE 1